OpenStudy (snowcrystal):

Can someone help me with this please? I will fan and medal. I dont understand this problem an im in a hurry.

OpenStudy (snowcrystal):

Graph the function and identify the domain and range y = -3x^2

OpenStudy (snowcrystal):

\[A.) domain: (-\infty , \infty) range : [0, \infty)\] \[B.) domain: (-\infty , \infty) range: (-\infty , 0]\] \[C.) domain: (-\infty, \infty) range: (-\infty, 0]\] \[D.) domain: (-\infty, \infty) range: [0,\infty)\]

OpenStudy (snowcrystal):

@MathLegend @hoblos

OpenStudy (snowcrystal):

I know some of the choices are the same but i cant put the pics of the graph up

OpenStudy (anonymous):